Grade 12th passGeneral Physics

A spring of constant K is cut into two parts b the ratio of length 1:2 The spring constant of larger one is?

Muskan Fatima
9 Years agoGrade 12th pass

1 Answer

Muhammad sajeer
9 Years ago
As we know,the product of spring constent and length will be a constent.ie,KL is content.K1L1=K2L2So,as per question, L1/L2=1/2=K2/k1So,k2=k1/2That means,spring content of long length string is half of spring constent of small. Hope u understand...